Find the function y=y(x) (for x>0) which satisfies the separable differential equation

dy/dx = (4+17x)/(xy^2). ;x>0


with the initial condition: y(1)=2
Answer #1

DATE TUDI fy day - se dane + faze da y3 = 4 lalalt 17 (x) te $ = 0 +27+0 ->C=-43 as they + 43 = 4 lalz/+ 27 (2) 3+43 = 3 [4 lDo give a thumbs up if you like this solution